A wealthy businessman opens a factory that manufactures a type of indoor soccer boot. He employs 16 workers each of whom is paid a monthly salary of R8000. It costs R140 to manufacture pair of boots and each pair sells for R300. Assume that no other costs are invloved, how many pairs of boots must be sold each month for the business to break even
Answer by mananth(16946)

Break even
sales = cost
let no of boots be x
workers 16
R8000 /month
wages = 16*8000 = 128000
140 to make a pair
140x
sales = 300x
300x= 140x+128000
160x=128000
/160
x=800 pairs